Hydrogen gas, iodine vapor and hydrogen iodide gas are added to an evacuated flask until the concentrations of H2 and I2 are both 2.0 M and that of HI equals 8.0M K=9.00 at 423C

The reaction vessel is heated to 675 where the equilibrium concentration of the HI is determined to be 6.0M What is the K value at 675C
Answer #1

Solution-
The reaction is written as

H2(g) + I2(g) <==> 2HI(g) --- Kc = 9.0

2M .......2M ............ 8M

Kc = [HI]^2 / [H2] [I2]

Q = (8)^2/ / (2)^2)

Q = 64 / 4

Q = 16

b)- Reaction H2cvt I 2 - 241 dorint -2. . Colpa E (3 (d) (r) 3 do 6 bK = 6 03) (3) = 4